# Fully Parallelized BP Decoding for Quantum LDPC Codes Can Outperform BP-OSD

### v2

@@
Abstract.
In this work, we propose a lightweight decoder based solely on belief-propagation (BP), augmented with a speculative post-processing strategy inspired by classical Chase decoding. Our method identifies unreliable bits via BP oscillation statistics, generates a set of modified test patterns, and decodes them in parallel using low-iteration BP. We demonstrate that our approach can achieve logical error rates comparable to or even better than BP-OSD, but has lower latency over its parallelization for a variety of bivariate bicycle codes, which significantly reduces decoding complexity.
1. Introduction
@@
Quantum low-density parity-check (qLDPC) codes have attracted significant attention in recent years due to their potential to encode more logical qubits than surface codes while maintaining a high threshold (Tillich and Zémor, 2014; Panteleev and Kalachev, 2021; Bravyi et al., 2024). However, decoding qLDPC codes remains a major challenge. To ensure reliable operation, especially in fault-tolerant quantum memory and computation, errors must be corrected both quickly and accurately, placing stringent demands on the decoding algorithm in terms of performance and computational efficiency.
@@
Belief propagation (BP)-based decoders, widely used in classical LDPC codes, are appealing due to their low complexity, parallelizability, and near-optimal performance. However, their effectiveness diminishes significantly when applied to qLDPC codes. This degradation is primarily due to inherent properties of qLDPC codes, such as degeneracy, the presence of many low-weight stabilizers, and the prevalence of trapping sets (Raveendran and Vasić, 2021), which hinder convergence and reliability of BP decoding.
@@
Many works have sought to address the limitations of BP-based decoders on qLDPC codes. In (Poulin and Chung, 2008), Poulin and Chung proposed several techniques to enhance convergence, including random freezing of variable nodes, perturbing priors, and colliding unsatisfied check nodes. In (Chytas et al., 2024), instead of performing static trapping set analysis prior to decoding, the authors identify bits affected by trapping sets dynamically during the decoding process. Once identified, the priors of these bits are modified, similar to the approach in (Poulin and Chung, 2008), to help the decoder escape local minima and converge. Raveendran et al. (Raveendran and Vasić, 2021) analyzed different types of trapping sets and proposed using a layered BP decoder to mitigate the effect of symmetric trapping sets. While this approach can reduce complexity, it often comes at the cost of increased decoding latency. More recently, Yin et al. (Yin et al., 2024) leveraged the degeneracy property of quantum LDPC codes to enhance BP decoding. By analyzing bit-wise marginal probabilities from BP, they selectively split rows in the parity-check matrix and modify the corresponding Tanner graph. This symmetry-breaking technique helps the decoder avoid convergence stalls caused by structural degeneracies in the code.
@@
Apart from efforts aimed at improving the BP decoder itself, several works focus on post-processing techniques to enhance decoding performance. A widely used approach is belief propagation combined with ordered statistics decoding (BP-OSD) (Panteleev and Kalachev, 2021; Roffe et al., 2020). While BP-OSD significantly enhances error correction performance, its reliance on a Gaussian elimination step during the OSD phase introduces substantial computational overhead. Specifically, this step incurs a complexity of $O(N^{3})$, where $N$ denotes either the code length (in the code-capacity error model) or the number of error mechanisms (in the circuit-level noise model). In contrast, each iteration of BP has only $O(N)$ complexity. This difference makes BP-OSD computationally expensive and less suitable for large-scale or real-time decoding applications. To address this limitation, recent works such as (Wolanski and Barber, 2024; Hillmann et al., 2024) propose partitioning the Tanner graph into several clusters. This localized decoding approach reduces the size of the matrices involved in Gaussian elimination, thereby lowering the overall computational burden without substantially compromising decoding performance.
@@
In this work, rather than aiming to improve BP itself, we focus on achieving decoding performance comparable to BP-OSD while avoiding computationally expensive post-processing steps such as Gaussian elimination. To this end, we propose a lightweight post-processing technique based entirely on BP and demonstrate that, with appropriate design and a speculative decoding strategy, it can closely match the logical error rates achieved by BP-OSD. Through simulations on various bivariate bicycle codes, we show that the proposed decoder performs similarly in terms of syndrome analysis to BP-OSD with a combination-sweep of order 10. Additionally, the proposed BP decoder requires lower latency given it is fully parallelizable, which means it can outperform BP-OSD in execution efficiency, making it a promising candidate for scalable and efficient decoding of quantum LDPC codes in practical fault-tolerant quantum computing systems.
Throughout the paper, we adopt the widely used min-sum variant of BP because of its simplicity and computational efficiency. Nonetheless, our approach could potentially benefit from incorporating more advanced BP-based techniques.
2. Background
2.1. Quantum LDPC Codes
@@
(1) $\mathcal{C}=\{\ket{\psi}|\>s\ket{\psi}=\ket{\psi},\>\forall s\in\mathcal{S}\}.$
An $\llbracket n,k,d\rrbracket$ stabilizer code can be defined by $n-k$ independent stabilizers, allowing us to encode $k$ qubits of logical information into an $n$-qubit block tolerating up to ${\lfloor(d-1)/2\rfloor}$ errors. CSS codes are an important class of stabilizer with two sets of stabilizers, $X$-type and $Z$-type, represented by parity-check matrices $H_{X}$ and $H_{Z}$, respectively. Each row in a parity-check matrix corresponds to a stabilizer generator, and each column corresponds to a physical qubit. A “1” entry indicates an $X$ or $Z$ operator (depending on whether it is in $H_{X}$ or $H_{Z}$), while a “0” indicates the identity. Consequently, an $X$-type stabilizer acts as $X$ or the identity on each qubit, and a $Z$-type stabilizer acts as $Z$ or the identity on each qubit. Errors can therefore be corrected by handling $Z$ errors and $X$ errors separately. Since all stabilizers must commute with each other, it follows directly that for a CSS code $H_{X}H_{Z}^{T}=0$.
2.2. Decoding Problem
@@
(3) $[\hat{E}]=\arg\max_{[E]:\text{synd}(E)=\bm{s}}\Pr([E]),$
where $[E]$ denotes the equivalence class of errors under the stabilizer group and $\Pr([E])$ is the total probability of all errors in that class.
While the classical decoding problem in Eq. (2) is NP-hard, its quantum counterpart in Eq. (3) is even more complex, being #P-complete. In practice, optimal decoding is computationally intractable, so efficient decoders typically aim to approximate the solution to Eq. (2) with good performance under realistic noise models.
As CSS codes can be decoded separately on the $X$- and $Z$-error bases, each decoding problem can be treated as a classical decoding task and addressed using classical decoding algorithms. For example, BP, the most commonly used decoder for classical LDPC codes, is also widely applied in qLDPC codes. Given an $M\times N$ parity-check matrix $H$, let $v_{1},\dots,v_{N}$ denote the variable nodes (corresponding to the columns of $H$) and $c_{1},\dots,c_{M}$ the check nodes (corresponding to the rows). The normalized min-sum algorithm, a widely used variant of BP can be described as follows:
  1. (1)
@@
    (5) $l_{v_{i}\rightarrow c_{j}}=l^{ch}_{v_{i}}+\sum_{c_{j^{\prime}}\in N(v_{i})\setminus\{c_{j}\}}l_{c_{j^{\prime}}\rightarrow v_{i}}\>\>,$
@@
    (6) $l_{c_{j}\rightarrow v_{i}}=(-1)^{s_{j}}\cdot\alpha\min_{v_{i^{\prime}}\in N(c_{j})\setminus\{v_{i}\}}|l_{v_{i^{\prime}}\rightarrow c_{j}}|\cdot\prod_{v_{i^{\prime}}\in N(c_{j})\setminus\{v_{i}\}}\text{sign}(l_{v_{i^{\prime}}\rightarrow c_{j}}),$
@@
    (7) $l^{out}_{v_{i}}=l^{ch}_{v_{i}}+\sum_{c_{j^{\prime}}\in N(v_{i})}l_{c_{j^{\prime}}\rightarrow v_{i}}\>\>,$
@@
    (8) $\hat{e_{i}}=\begin{cases}0&\text{if}\quad l_{v_{i}}^{out}>0\\ 1&\text{otherwise}\end{cases},$
@@
In the classical setting, BP performs close to optimal. However, for qLDPCs, the performance of BP degrades significantly: The decoder may fail to converge due to the high degeneracy of quantum codes and the absence of soft information for each bit, which is typically available in classical scenarios.
3. BP Behavior Analysis: A Case Study
@@
In this section, we analyze the behavior of BP-based decoders on qLDPC codes using the $\llbracket 144,12,12\rrbracket$ “gross” code from (Bravyi et al., 2024) as a representative case study. While specific to this code, the analysis provides general insights into the behavior of BP decoding on qLDPC codes and offers guidance for improving decoder performance. The BP decoder used below is a min-sum decoder with adaptive damping factor $\alpha=1-2^{i}$, where $i$ is the current number of iterations.
3.1. Number of Iterations
@@
Figure 1 shows the non-convergence rate of the min-sum decoder under the circuit-level noise model, where $p$ denotes the physical error rate. The curves are obtained by simulating 10,000 samples for each of $p=0.001$ and $p=0.002$, both representative values below the decoding threshold. For each sample, we record the number of iterations required for convergence and compute the cumulative distribution. The non-convergence rate at iteration $i$ is defined as the fraction of samples that have not converged within $i$ iterations—that is, $1$ minus the cumulative convergence rate.
As Fig. 1 indicates, in most cases, BP converges within a small number of iterations. For instance, at $p=0.001$, the average number of iterations is merely 8.9. Even at higher error rates, such as $p=0.002$, the average number of iterations remains low, although the tail becomes longer. Notably, cases that do not converge within the early iterations rarely benefit from increasing the iteration count further. This observation motivates an alternative strategy: Rather than extending the number of BP iterations, we can vary the inputs to the BP decoder while keeping the maximum number of iterations small. If these varied inputs have better independent chances of successful decoding, then running multiple instances in parallel allows us to exponentially suppress the logical error rate without incurring significant decoding latency.
@@
Figure 1. Ratio of unsuccessful BP decoding ($1-$convergence rate) on the $\llbracket 144,12,12\rrbracket$ code under the circuit-level noise model. The maximum number of decoder iterations is set to 1,000 and number of samples is 10,000.
3.2. Oscillation
Figure 2. Precision and recall probabilities of candidate bit selection on the $\llbracket 144,12,12\rrbracket$ code. We evaluate the correlation between candidate bits and actual error locations by identifying the top 50 most frequently flipped bits among approximately 8,000 error mechanisms. The decoder is run with a maximum of 50 iterations, and statistics are collected over 1,000 decoding failures. This analysis reveals how well bit-level oscillation can serve as a heuristic for error localization.

### v3

@@
Abstract
This work presents a hardware-efficient and fully parallelizable decoder for quantum LDPC codes that leverages belief propagation (BP) with a speculative post-processing strategy inspired by classical Chase decoding algorithm. By monitoring bit-level oscillation patterns during BP, our method identifies unreliable bits and generates multiple candidate vectors to selectively flip syndromes. Each modified syndrome is then decoded independently using short-depth BP, a process we refer to as BP-SF (syndrome flip). This design eliminates the need for costly Gaussian elimination used in the current BP-OSD approaches. Our implementation achieves logical error rates comparable to or better than BP-OSD while offering significantly lower latency due to its high degree of parallelism for a variety of bivariate bicycle codes. Evaluation on the $\llbracket 144,12,12\rrbracket$ bivariate bicycle code shows that the proposed decoder reduces average latency to approximately $70\%$ of BP-OSD. When post-processing is parallelized the average latency is reduced by $55\%$ compared to the single process implementation, with the maximum latency reaching as low as $18\%$. These advantages make it particularly well-suited for real-time and resource-constrained quantum error correction systems.
footnotetext: This paper has been accepted to and published at the IEEE International Symposium on High-Performance Computer Architecture (HPCA) 2026.
I Introduction
@@
One of the central challenges in quantum computing is achieving fault-tolerant quantum computation (FTQC), which requires the use of quantum error correction (QEC) codes to correct errors arising from noisy quantum devices. Significant efforts [7, 31, 30, 6, 29] have been made toward designing decoders and hardware architectures for various QEC codes. Among the QEC codes, quantum low-density parity-check (qLDPC) codes have attracted significant attention in recent years due to their potential to encode more logical qubits than surface codes while maintaining a high threshold [27, 18, 2]. However, decoding qLDPC codes remains a major challenge. To ensure reliable operation, especially in fault-tolerant quantum memory and computation, errors must be corrected both quickly and accurately. In practice, decoders must keep pace with the rate of syndrome extraction to prevent data backlog [25], placing stringent demands on both the decoding algorithms and their implementations in terms of performance and computational efficiency.
@@
Belief propagation (BP)-based decoders, widely used in classical LDPC codes, are appealing due to their low complexity, parallelizability, and near-optimal performance [21, 8, 24, 37]. However, their effectiveness diminishes significantly when applied to qLDPC codes. This degradation is primarily due to inherent properties of qLDPC codes, such as degeneracy, the presence of many low-weight stabilizers, and the prevalence of trapping sets, which hinder convergence and reliability of BP decoding [20].
@@
Many works have sought to address the limitations of BP-based decoders on qLDPC codes. In [19], Poulin and Chung proposed several techniques to enhance convergence, including random freezing of variable nodes, perturbing prior information, and colliding unsatisfied check nodes. Similar approaches are also adopted in [35] to freeze a node based on posterior information. In [5], instead of performing static trapping set analysis prior to decoding, Chytas et al. identified oscillation bits affected by trapping sets dynamically during the decoding process. Once identified, the posterior information of these bits is modified, similar to the approach in [19], to help the decoder escape local minima and converge. Raveendran et al. [20] analyzed different types of trapping sets and proposed using a layered BP decoder to mitigate the effect of symmetric trapping sets. While layered decoder can reduce complexity, it often comes at the cost of increased decoding latency as layered decoders are serial. More recently, Yin et al. [36] leveraged the degeneracy property of quantum LDPC codes to enhance BP decoding. By analyzing bit-wise marginal probabilities from BP, they selectively split rows in the parity-check matrix and modify the corresponding Tanner graph. This symmetry-breaking technique helps the decoder avoid convergence stalls caused by structural degeneracies in the code. In general, these methods typically involve modifications to the graph structure, prior, or posterior information in BP decoding to constrain the decoder, which effectively reduce the search space to facilitate convergence. In [11], Gong et al. proposed guided decimation guessing (GDG), a method based on tracking the decoding history of BP to accelerate its convergence. They also employ a tree-search-like strategy that keeps multiple decimation paths open to correct errors as a decoding ensemble. However, the decision tree-like structure of the “guessing” phase restricts the algorithm’s potential for further parallelization. During the preparation of this work, Müller et al. [17] proposed Relay-BP, a hardware-friendly and real-time decoder constructed by chaining multiple Mem-BP [4] decoders with varying memory strengths. While this approach achieves a logical error rate significantly lower than the BP-OSD decoder, its sequential chaining structure imposes a latency overhead that prevents the different decoding stages from being parallelized.
Figure 1: (a) Example of BP failing to converge due to oscillations. Red Xs denote bits identified as erroneous by BP, and yellow squares represent unsatisfied syndrome checks. All four bits are oscillating. (b) One oscillating bit (e.g., the rightmost one) is selected, and its neighboring syndromes are flipped. BP then converges since the two competing error patterns now have different weights. (c) After convergence, the selected bit is flipped back to restore consistency with the original input syndrome.
@@
Apart from efforts aimed at improving the BP decoder itself, several works focus on post-processing techniques to enhance decoding performance. A widely used approach is belief propagation combined with ordered statistics decoding (BP-OSD) [18, 22]. While BP-OSD significantly enhances error correction performance, its reliance on a Gaussian elimination step during the OSD phase introduces substantial computational overhead. Specifically, this step incurs a complexity of $O(N^{3})$ [18], where $N$ denotes either the code length (in the code-capacity error model) or the number of error mechanisms (in the circuit-level noise model). In contrast, each iteration of BP has a computational complexity of only $O(N)$ [21]. This difference makes BP-OSD computationally expensive and less suitable for large-scale or real-time decoding applications. To address this limitation, recent works such as [34, 13] proposed partitioning the Tanner graph into several clusters. This localized decoding approach reduces the size of the matrices involved in Gaussian elimination, thereby lowering the overall computational burden without substantially compromising decoding performance. These post-processing techniques provide a fallback when BP decoding fails, but they often involve complex data structures and control flows, making them difficult to implement efficiently in hardware.
In this work, rather than focusing on improving BP or OSD individually, we propose that we can achieve error rate performance comparable to BP-OSD while offering significantly lower latency. The key insights behind our approach, which we refer to as BP-SF, are twofold: (i) Most BP decoding failures are caused by a small subset of bits referred to as oscillating bits. By flipping some of these bits in the syndrome domain, BP can often converge rapidly. (ii) Due to the low complexity and inherently parallel nature of BP, we can speculatively perform multiple decoding attempts in parallel, incurring minimal additional latency. Our technique can also be viewed as analogous to providing different starting points to the optimizer, helping it escape local minima and find a global one.
@@
We give a brief visualization in Figure 1 to illustrate our purely BP-based algorithm. We demonstrate that, with appropriate implementation and design, it can closely match the logical error rates achieved by BP-OSD. Through simulations on a variety of quantum LDPC codes, including the $\llbracket 72,12,6\rrbracket$, $\llbracket 144,12,12\rrbracket$, and $\llbracket 288,12,18\rrbracket$ bivariate bicycle codes; the $\llbracket 126,12,10\rrbracket$ and $\llbracket 154,6,16\rrbracket$ coprime bivariate bicycle codes; the $\llbracket 225,16,8\rrbracket$ subsystem hypergraph product simplex code and the $\llbracket 254,28\rrbracket$ generalized bicycle code, we show that the proposed decoder performs similarly to BP-OSD with a combination-sweep of order 10. Additionally, the proposed BP decoder requires lower latency as it is fully parallelizable, which means it can outperform BP-OSD in execution efficiency, making it a promising candidate for scalable and efficient decoding of quantum LDPC codes in practical fault-tolerant quantum computing systems.
Contributions:
  • BP Decoding Analysis: We analyze the behavior of the BP decoding algorithm using the $\llbracket 144,12,12\rrbracket$ code as a case study, revealing a long-tail distribution in the number of iterations and a strong correlation between oscillating bits and actual error locations.
  • Oscillation-Guided Speculative Decoding: We propose a fully parallelizable post-processing technique that leverages bit-level oscillation during BP decoding to identify unreliable bits and generate test vectors, enabling Chase-like decoding without requiring Gaussian elimination.
  • Efficient Implementation: We implement the proposed decoder in both serial and multi-process CPU versions, achieving faster runtime than BP-OSD while maintaining comparable logical error rates across various qLDPC codes. We also explore GPU implementation potential, and provide a pessimistic upper bound for GPU decoding time.
II Background
II-A Quantum LDPC Codes
@@
[…]
An $\llbracket n,k,d\rrbracket$ stabilizer code can be defined by nkn-k independent stabilizers, allowing us to encode $k$ qubits of logical information into an $n$-qubit block tolerating up to $\lfloor(d-1)/2\rfloor$} errors. CSS codes are an important class of stabilizer with two sets of stabilizers, $X$-type and $Z$-type, represented by parity-check matrices $H_{X}$ and $H_{Z}$, respectively. Each row in a parity-check matrix corresponds to a stabilizer generator, and each column corresponds to a physical qubit. A “1” entry indicates an $X$ or $Z$ operator (depending on whether it is in $H_{X}$ or $H_{Z}$), while a “0” indicates the identity. Consequently, an $X$-type stabilizer acts as $X$ or the identity on each qubit, and a $Z$-type stabilizer acts as $Z$ or the identity on each qubit. Errors can therefore be corrected by handling $Z$ errors and $X$ errors separately. Since all stabilizers must commute with each other, it follows directly that for a CSS code $H_{X}H_{Z}^{T}=0$. If both $H_{X}$ and $H_{Z}$ are sparse matrices, the code is a CSS-type qLDPC code. The sparsity of these matrices offers a key advantage: syndrome extraction can be performed using fewer quantum gates, thereby reducing circuit depth and potential error accumulation. In this sense, the surface code can be regarded as a special case of qLDPC codes, characterized by strictly local, nearest-neighbor interactions on a 2D lattice. More general qLDPC codes, in contrast, typically exhibit higher connectivity and non-local parity-check relationships. This allows them to encode more logical qubits while maintaining a large code distance, but it also makes matching-based algorithms (commonly used for surface codes) less effective, since the increased connectivity introduces hyperedges into the decoding graph.
II-B Decoding Problem
@@
$\hat{E}]=\arg\max_{[E]:\text{synd}(E)=\bm{s}}\Pr([E]),$ (3)
where $[E]$ denotes the equivalence class of errors under the stabilizer group and $\Pr([E])$ is the total probability of all errors in that class. While the classical decoding problem in Eq. (2) is NP-hard [1], its quantum counterpart in Eq. (3) is even more complex, being #P-complete [14]. In practice, optimal decoding is computationally intractable, so efficient decoders typically aim to approximate the solution to Eq. (2) with good performance under realistic noise models.
As CSS codes can be decoded separately on the $X$- and $Z$-error bases, each decoding problem can be treated as a classical decoding task and addressed using classical decoding algorithms. For example, BP, the most commonly used decoder for classical LDPC codes, is also widely applied in qLDPC codes. This is because the sparsity of qLDPC and LDPC codes can be effectively exploited by BP-based decoders, which rely on the assumption of independent probability updates. But this assumption is only valid when the parity-check matrix is sparse. As the matrix becomes denser, the variable nodes exhibit stronger correlations, violating the independence assumption. Consequently, although the BP decoder can, in principle, be applied to denser codes, its empirical performance tends to degrade due to the increased dependency within the graph structure. Given an $M\times N$ parity-check matrix $H$, let $v_{1},\dots,v_{N}$ denote the variable nodes (corresponding to the columns of $H$) and $c_{1},\dots,c_{M}$ the check nodes (corresponding to the rows). The normalized min-sum algorithm, a widely used variant of BP can be described as follows:
  1. 1.
@@
    $l_{v_{i}\rightarrow c_{j}}=l^{ch}_{v_{i}}+\sum_{c_{j^{\prime}}\in N(v_{i})\setminus\{c_{j}\}}l_{c_{j^{\prime}}\rightarrow v_{i}}\>\>,$ (5)
@@
    $\begin{split}l_{c_{j}\rightarrow v_{i}}=&(-1)^{s_{j}}\cdot\alpha\min_{v_{i^{\prime}}\in N(c_{j})\setminus\{v_{i}\}}|l_{v_{i^{\prime}}\rightarrow c_{j}}|\cdot\\ &\prod_{v_{i^{\prime}}\in N(c_{j})\setminus\{v_{i}\}}\text{sign}(l_{v_{i^{\prime}}\rightarrow c_{j}}),\end{split}$ (6)
@@
    $l^{out}_{v_{i}}=l^{ch}_{v_{i}}+\sum_{c_{j^{\prime}}\in N(v_{i})}l_{c_{j^{\prime}}\rightarrow v_{i}}\>\>,$ (7)
@@
    $\hat{e_{i}}=\begin{cases}0&\text{if}\quad l_{v_{i}}^{out}>0\\ 1&\text{otherwise}\end{cases},$ (8)
@@
III BP Behavior Analysis: A Case Study
@@
In this section, we analyze the behavior of BP-based decoders on qLDPC codes using the $\llbracket 144,12,12\rrbracket$ “gross” code from [2] as a representative case study. While specific to this code, the analysis provides general insights into the behavior of BP decoding on qLDPC codes and offers guidance for improving decoder performance. The BP decoder used below is a min-sum decoder as described in Eq. (6) with an adaptive damping factor of $\alpha=1-2^{i}$, where $i$ is the current number of iterations.
III-A Number of Iterations
@@
Figure 2 shows the ratio of syndromes failed to converge under the circuit-level noise model, where $p$ denotes the physical error rate. The curves are obtained by simulating 10,000 samples for each of $p=0.001$ and $p=0.002$, both representative values below the threshold. For each sample, we record the number of iterations required for convergence and compute the cumulative distribution. This non-convergence rate at iteration $i$ is defined as the fraction of samples that have not converged within $i$ iterations—that is, $1$ minus the cumulative convergence rate.
@@
Figure 2: Ratio of unsuccessful BP decoding (11-convergence rate) on the $\llbracket 144,12,12\rrbracket$ code under the circuit-level noise model. The maximum number of decoder iterations is set to 1,000 and number of samples is 10,000.
As Fig. 2 indicates, in most cases, BP converges within a small number of iterations. For instance, at $p=0.001$, the average number of iterations is merely 8.9 despite setting maximum number of iterations to 1,000. Even at higher error rates, such as $p=0.002$, the average number of iterations remains low, although the tail becomes longer. Notably, cases that do not converge within the early iterations rarely benefit from increasing the iteration count further. This observation motivates an alternative strategy: Rather than extending the number of BP iterations, we can vary the inputs to the BP decoder while keeping the maximum number of iterations small. If these varied inputs have better independent chances of successful decoding, then running multiple instances in parallel allows us to exponentially suppress the logical error rate without incurring significant decoding latency.
III-B Oscillation